# Sky localization of gravitational waves from eccentric binaries

### v1

@@ V Conclusions
Through simulations, we have shown that the orbital eccentricity in inspiraling neutron stars can be used to improve their sky localization from gravitational wave observations. However, this necessitates that existing algorithms be optimized for detecting and localizing eccentric binaries. Here we have demonstrated a semi-Bayesian approach to conduct such localizations that uses the output of a matched-filter eccentric search. While we have discussed the implications for binary neutron stars, this approach can be easily extended to any modeled eccentric binary merger. The chances of observations also crtitically depend on the astrophysical abundancy of the sources. Even if few detectable sources exist within the reach of the detectors, current analyses can miss substantial SNRs and hence, the opportunity to pin down their sky positions. With improved low-frequency sensitivities, lighter binaries with greater early-inspiral cycles are relatively more prone to losing signal power as shown in Fig. 8. We have explored qualitative improvements with an eccentric early-warning system over a non-eccentric one with the current sensitivities of the detectors. We have identified some important cases with eccentric early-warning localizations from the EM-observations viewpoint.
@@
While we have used only a two detector HL network for demonstration, it is useful to investigate the cases when more detectors are included such as Virgo and KAGRA. Given the semi-Bayesian nature of our approach, the localization method described here is expected to readily handle sub-threshold signals in any given detector(s). Additional detectors should guide the network’s senstivity to the detectable binaries. Further improvements in localizing the astrophysical sources are awaited when LIGO India comes online (Saleem et al., 2021).
@@
We have conducted studies on simulated data and provided only rough latency estimates without considering practical overheads. However, end-to-end tests of the capabilties of the eccentric early-warning system requires deployment of the search with a real-time interface. Fortunately, such an infrastructure is already operational within the LIGO-Virgo Collaboration (Magee et al., 2021; Chaudhary et al., 2024). Since an eccentric EW search can accumulate greater SNRs in general, they can potentially reach a given significance threshold quicker, allowing greater EW times for eccentric sources. With the improved low-frequency senstivities of the instruments, we find that a dedicated eccentric EW system is desirable as shown in Fig. 9. Such system will improve the chances for the first and the subsequent EW observations. Though the localization algorithm presented in this work implements a computationally effective strategy, further optimizations are needed to realize true EW observations. Alternately, current algorithms can be optimized to include eccentricity as an input parameter. Machine learning based approaches can already reach very low latencies (Chatterjee and Wen, 2023).
Figure 11: Self-consistency of the FAR estimates with varying coincidence windows investigated for the EW searches using datasets from O3b. The gray region represents the 95% Poisson uncertainty estimates around the expected line. The largest coincidence window chosen till including fairly all injections shown in Fig. 7. Time-slide interval is twice the coincidence window.
@@ VI Acknowledgments
We thank K. Rajesh Nayak for the providing valuable feedback on this work. We have used ligo.skymap, PyCBC, dynesty, numpy, cupy, and other packages for performing the analyses. Some of the results in this paper have been derived using the healpy and HEALPix package. The research and development work was carried out at the Center of Excellence in Space Sciences India (CESSI). CESSI is a multi-institutional Center of Excellence hosted by the Indian Institute of Science Education and Research (IISER) Kolkata and has been established through funding from the Ministry of Education, Government of India. We acknowledge the use of the Sarathi computing cluster at IUCAA for computational/numerical work. A portion of the GPU benchmarking was performed on the Kepler cluster hosted by the Department of Physical Sciences, IISER Kolkata. Partial support for the project was received from the Council for Scientific and Industrial Research (CSIR), India through File No:09/921(0272)/2019-EMR-I. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The article has a LIGO Document number LIGO-P2500001. We thank Gaurav Waratkar for carefully reading the manuscript.
@@ Appendix A Fallback localizations and the uncertainty estimates
In the absence of eccentric sources, the algorithm retreats to the localization accuracies optimized for the aligned-spin case. To demonstrate, we generate 1000 injections from the same population as given in Table. 1 but with $\varepsilon=0$. These are recovered with the aligned-spin PSO search. The recovered injections are localized using BAYESTAR and the localization algorithm introduced in this work. These is summaried in Fig. 10. The correction factors needed to obtain the pp-plots are obtained empirically. Such factors are used to rescale the log-likelihood ratio while generating the posterior samples. We have generally observed that within a given method, the factor shifts more from unity when there is more deviation of the estimated parameters from the injected ones. This aspect has been discussed in detail in (Duverne et al., 2024). Note that the tunable parameter can be modified depending on the target population of sources and does not affect the accuracy of localizations. We do not expect these values to change considerably when the algorithm is used with real signals with the reasonable assumption that real instrument noise can be treated as approximately Gaussian except for the occasional instances of glitches localized in time.
Appendix B Backgrounds event rates in early-warning searches
@@
In this section, we explore the suitability of increasing the time-coincidence window for early-warning searches beyond the standard value used in the full-bandwidth searches. A coincidence window of 15 ms for the HL detector network can potentially miss several astrophysical EW candidates due to timing uncertainties of EW triggers, though, such events are nominally expected to feature in an FB analysis. Thus, to improve the sensitivities of the EW analyses, a large coincidence window should be chosen. For the simulated population of the sources, the maximum of such windows is expected to recover the signals with coincidence. However, a larger coincidence window is also expected to allow more triggers from non-astrophysical origin to enter into coincidence. However, increasing the width of the coincidence window also increases the chances that triggers of noise origin form undesirable coincidences. To investigate such chances for the PSO-based searches, we generate EW triggers in O3b data (Abbott et al., 2021, 2023). Here we use datasets from LIGO Hanford and LIGO Livingston to generate EW triggers from the same parameter space described earlier in this work. To compute the background events, the triggers from the individual detectors are time-slided with respect to each other. These are described in (Pal and Nayak, 2024). Here, the time-slide interval is assumed to be twice as large as a time-coincidence window used. Thus, we repeat this procedure with the same triggers but by varying width of time-window beyond 15 ms for the HL network. The maximum value used is motivated from the possible timing errors associated with a given frequency band estimated in Fig. 7. To accomodate triggers passing a smaller SNR threshold, a larger width for the coincidence window may be necessary. We observe that such triggers result into foreground coincidences with self-consistent inverse false alarm rate (IFAR) estimates. Thus, we broadly conclude that allowing for larger coincidence window to accomodate astrophysical signals does not interfere with their significance calculation.

### v4

@@ V Conclusions
Through simulations, we have shown that the orbital eccentricity in inspiraling neutron stars can be used to improve their sky localization from gravitational wave observations. However, this necessitates that existing algorithms be optimized for detecting and localizing eccentric binaries. Here we have demonstrated a semi-Bayesian approach to conduct such localizations that uses the output of a matched-filter eccentric search. While we have discussed the implications for binary neutron stars, this approach can be easily extended to any modeled eccentric binary merger. The chances of observations also crtitically depend on the astrophysical abundancy of the sources. Even if few detectable sources exist within the reach of the detectors, current analyses can miss substantial SNRs and hence, the opportunity to pin down their sky positions. With improved low-frequency sensitivities, lighter binaries with greater early-inspiral cycles are relatively more prone to losing signal power. We have explored qualitative improvements with an eccentric early-warning system over a non-eccentric one with the current sensitivities of the detectors. We have identified some important cases with eccentric early-warning localizations from the EM-observations viewpoint.
@@
While we have used only a two detector HL network for demonstration, it is useful to investigate the cases when more detectors are included such as Virgo and KAGRA. Given the semi-Bayesian nature of our approach, the localization method described here is expected to readily handle sub-threshold signals in any given detector(s). Additional detectors should guide the network’s senstivity to the detectable binaries. Further improvements in localizing the astrophysical sources are awaited when LIGO India comes online (Saleem et al., 2021; Shukla et al., 2024).
@@
We have conducted studies on simulated data and provided only rough latency estimates without considering practical overheads. However, end-to-end tests of the capabilties of the eccentric early-warning system requires deployment of the search with a real-time interface. Fortunately, such an infrastructure is already operational within the LIGO-Virgo Collaboration (Magee et al., 2021; Chaudhary et al., 2024). Since an eccentric EW search can accumulate greater SNRs in general, they can potentially reach a given significance threshold quicker, allowing greater EW times for eccentric sources. With the improved low-frequency senstivities of the instruments, we find that a dedicated eccentric EW system is desirable as discussed further in Appendix D. Such a system will improve the chances for the first and the subsequent EW observations. Though the localization algorithm presented in this work implements a computationally effective strategy, further optimizations are needed to practically realize EW observations. Alternately, current algorithms can be optimized to include eccentricity as an input parameter. Machine learning based approaches can already reach very low latencies (Chatterjee and Wen, 2023).
Configuration Aligned-spin injections Spin-aligned eccentric injections
This work BAYESTAR This work BAYESTAR
Injected 0.9 0.5
Estimated 0.55 0.8 0.5 0.7
Table 3: The log-likelihood ratio (or the SNR) rescaling factors used to obtain the pp-plots in Fig. 4 and Fig. 13.
@@ VI Acknowledgments
We thank K. Rajesh Nayak for the providing valuable feedback on this work. We have used ligo.skymap, PyCBC, dynesty, numpy, cupy, and other packages for performing the analyses. Some of the results in this paper have been derived using the healpy and HEALPix package. The research and development work was carried out at the Center of Excellence in Space Sciences India (CESSI). CESSI is a multi-institutional Center of Excellence hosted by the Indian Institute of Science Education and Research (IISER) Kolkata and has been established through funding from the Ministry of Education, Government of India. We acknowledge the use of the Sarathi computing cluster at IUCAA for computational/numerical work. A portion of the GPU benchmarking was performed on the Kepler cluster hosted by the Department of Physical Sciences, IISER Kolkata. Partial support for the project was received from the Council for Scientific and Industrial Research (CSIR), India through File No:09/921(0272)/2019-EMR-I. This material is based upon work supported by NSF’s LIGO Laboratory which is a major facility fully funded by the National Science Foundation. The article has a LIGO Document number LIGO-P2500001. We thank Gaurav Waratkar for carefully reading the manuscript. We sincerely acknowledge the valuable feedback from the anonymous referee.
@@ Appendix A Fallback localizations and the uncertainty estimates
In the absence of eccentric sources, the algorithm retreats to the localization accuracies optimized for the aligned-spin case. To demonstrate, we generate 1000 injections from the same population as given in Table. 1 but with $\varepsilon=0$. These are recovered with the aligned-spin PSO search. The recovered injections are localized using BAYESTAR and the localization algorithm introduced in this work. These are summaried in Fig. 13. The correction factors needed to obtain the pp-plots are obtained empirically, by systematically varying the value of the factor and checking the resulting pp-plots. The value at which a plot is reasonably consistent with the diagonal is noted for each case, as shown in Table. 3. Such factors are used to rescale the log-likelihood ratio while generating the posterior samples. We have generally observed that within a given method, the factor shifts more from unity when there is more deviation of the estimated parameters from the injected ones. This aspect has been discussed in detail in (Duverne et al., 2024). Note that the tunable parameter can be modified depending on the target population of sources and does not affect the accuracy of localizations. We do not expect these values to change considerably when the algorithm is used with real signals with the reasonable assumption that real instrument noise can be treated as approximately Gaussian except for the occasional instances of glitches localized in time.
Appendix B Distance estimates for eccentric injections
Figure 14: Summary of the distance estimates marginalized over the sky obtained under two different prior choices. We use a uniform prior for the luminosity distance (left) and a uniform prior in volume (right), determined by the sensitivity of the detectors, as described in the text. For these cases, a comparison of the mean errors (top) and the standard deviations (bottom) of the estimates is shown for the commonly found injections obtained with and without including eccentricity. The values in the parentheses indicate the average fractional error or the average standard deviation calculated over the localized sources.
Figure 15: Self-consistency of the FAR estimates with varying coincidence windows investigated for the EW searches using datasets from O3b. The gray regions represent the $1$, $2$, and $3\sigma$ Poisson uncertainty estimates around the expected line. The largest coincidence window chosen till including fairly all injections shown in Fig. 12. Time-slide interval is twice the coincidence window.
Figure 16: Average of the early-warning SNRs lost by various searches for the commonly recovered injections with O5-projected HL sensitivities as a function of the early-frequency bands.
In a 3D localization, the reconstructed distance to a source is a function of the sky direction, i.e, there exists a unique posterior probability distribution for the distance along any given sky direction, in general. To assess the distance estimate, the distribution along the direction corresponding to the most probable pixel of a skymap can be used. However, the true sky position may not be be along a given inferred direction, so we use the distribution marginalized over all sky directions. The mean and the standard deviation of the distribution can be extracted from the header (DISTMEAN and DISTSTD) of the 3D HEALPix maps. These are described in (Singer et al., 2016a, b). Here we discuss the estimates obtained with the eccentric and the non-eccentric analyses under two different priors choices, as shown in Fig. 14. We observe that, on average, the errors in the estimates are smaller for the eccentric case than for the non-eccentric one. Further, the standard deviation of the estimates of the localized sources is generally smaller in the eccentric case. These are consistent across both the prior choices- a uniform prior in the distance itself and a uniform prior in the volume, each limited by a fiducial maximum reach of the detectors approximately determined by the simulated sensitivity. However, the errors are generally smaller for the uniform prior in distance that we attribute to the choice of the injection distribution, which is also uniform in distance. The use of astrophysically informed priors can be further explored in the future. We obtain negligible impact on the accuracies of the areal localizations due to the choice of these priors as long as the priors cover the injection space.
Appendix C Backgrounds event rates in early-warning searches
@@
In this section, we explore the suitability of increasing the time-coincidence window for early-warning searches beyond the standard value used in the full-bandwidth searches. A coincidence window of 15 ms for the HL detector network can potentially miss several astrophysical EW candidates due to timing uncertainties of EW triggers, though, such events are expected to feature in an FB analysis. Thus, to improve the sensitivities of the EW analyses, a large coincidence window may be chosen. However, a larger coincidence window is also expected to allow more triggers of non-astrophysical origin to enter into coincidence. This would result into undesirable events, increasing false alarms. To investigate such chances for the PSO-based searches, we generate EW triggers in O3b data (Abbott et al., 2021, 2023). Here we use datasets from LIGO Hanford and LIGO Livingston to generate EW triggers from the same parameter space described earlier in this work. To compute the background events, the triggers from the individual detectors are time-slided with respect to each other. These processes are described in (Pal and Nayak, 2024). Here, the time-slide interval is assumed to be twice as large as a time-coincidence window used. We repeat this procedure with the same triggers but by varying width of time-window beyond 15 ms for the HL network. The maximum value used is motivated from the possible timing errors associated with a given frequency band estimated in Fig. 12. To accomodate triggers passing a smaller SNR threshold, a larger width for the coincidence window may be necessary. We observe that such triggers result into foreground coincidences with self-consistent inverse false alarm rate (IFAR) estimates, as shown in Fig. 15. Thus, we conclude that allowing for a larger coincidence window to accomodate a larger fraction of astrophysical signals does not interfere with their significance calculation in general.
Appendix D SNR loss in early-warning searches
Here we investigate the SNR loss from the PSO-based early-warning searches with several search-parameter combinations, in O5-like scenarios. The injection population is kept fixed as described earlier in Sec. IV. We use three combinations of the search parameters- (1) aligned-spins with eccentricity, (2) aligned-spins only, and (3) eccentricity only. The search configurations are otherwise kept same, including the template-sampling. For a similar discussion, refer to (Pal and Nayak, 2025). The average SNR losses for the commonly found injections as a function of the EW frequency ranges are plotted in Fig. 16. We observe that the SNR loss by the EW search with both aligned-spin and eccentricity outperforms other searches at all frequency ranges. We also find that at very low frequency ranges (notably, in the 10-20 Hz), the non-spinning but eccentric search recovers SNRs closely to the search with both aligned-spins and eccentricity. Note the non-eccentric but aligned-spin search suffers maximum and significant loss in the SNR in this frequency range. This indicates that while eccentricity is important to be included, the aligned-spin components could possibly have only sub-dominant effects at such low frequencies.